OpenStudy (anonymous):

How can I solve the equation -x^2 + 5x + 180?

OpenStudy (solomonzelman):

\huge\color{blue}{ \frac{±πθ}{2} } \(\Large\color{blue}{ \bf -x^{2}+5x+180=0 }\) multiply the equation times -1, then... \(\Large\color{blue}{ \bf ■~~quadratic~~formula: }\) \(\Large\color{blue}{ \bf x^{2}-5x-180=0 }\) \(\Large\color{blue}{ \bf \frac{-(-5)± \sqrt{(-5)^{2}-4(1)(-180)} }{2(1)} }\)

OpenStudy (anonymous):

Thank you!

OpenStudy (solomonzelman):

You welcome !